# Enhancing Agentic RL with Progressive Reward Shaping and Value-based Sampling Policy Optimization

### v1

@@ A Dynamic Advantage Weight Perspective of VSPO
VSPO can be interpreted as task-value-based reweighting rather than random sampling. Consider the general policy gradient (sutton1999policy) with advantage function $A_{t}$:
@@
To evaluate the effectiveness of PRS and VSPO in training LLM-based TIR agents, we conduct experiments on the following two types of long-horizon reasoning tasks: short-form QA and long-form QA. In these tasks, the agent has access to at least one information retrieval tool to answer the question.
In short-form QA, we use Qwen2.5-7B-Instruct and in long-form QA, we use the Qwen3-14B (qwen3technicalreport) as the training model. All training uses full parameter training.
We adopt Trinity-RFT (pan2025trinity) as our training framework.
For VSPO, we set $T=1$ and $\alpha=2$. The rollout num of VSPO, CISPO and GRPO is 5. The learning rate of all algorithms is 1e-6. Training is performed on a single node with 8 H100 GPUs. To optimize GPU memory usage, we enable gradient checkpointing and use Fully Sharded Data Parallel (FSDP) with CPU offloading.
@@
We manually annotate reference answers based on our internal knowledge base.
For expert data used in SFT, we use knowledge distillation from the Qwen3-235B-A22B (qwen3technicalreport) to construct high-quality trajectory data. Specifically, we collect online service logs from Qwen3-235B-A22B and apply rule-based filtering and LLM-as-a-Judge cleaning to produce expert demonstrations.
@@
Following (jin2025search), we use seven benchmark datasets: (1) General Question Answering: NQ , TriviaQA, and PopQA. (2) Multi-Hop Question Answering: HotpotQA, 2WikiMultiHopQA, Musique , and Bamboogle as Short-Form QA train and test data.
@@
We use Qwen3-235B-A22B-Instruct-2507 (qwen3technicalreport) as an LLM judge to evaluate responses against manually annotated reference answers. The evaluation prompts are provided in Appendix B.
@@
We use Exact Match (EM) as the evaluation metric, computed against the gold answers provided by each benchmark dataset.
@@
To verify the effectiveness of VSPO, we compare it with GRPO, PPO (schulman2017equivalence), CISPO (chen2025minimax) and SFT-only. In long-form QA, We also report the performance of Qwen3-235B-A22B as a reference for strong LLM capabilities.
For the reward design, we compare PRS with rule-based 0-1 reward to demonstrate the effectiveness of PRS. For short-form QA, using exact match as rule-based 0-1 reward (this is SEARCH-R1).
@@
Long-Form QA
The main results comparing VSPO with baseline methods across are presented in Table 1. VSPO surpasses most RL baselines (GRPO, PPO, CISPO) and the SFT-only setting on $Q_{simple}$, $Q_{multiq}$ and $Q_{multim}$. Compared to untrained, $Q_{simple}$, $Q_{multiq}$ and $Q_{multim}$ achieve 18.26% and 81.24% and 10% average relative improvement, which demonstrating that VSPO yields more effective policy updates in long-horizon tool-use scenarios.
$Q_{simple}$ $Q_{multiq}$ $Q_{multim}$
Qwen3-235B-A22B 0.8125 0.600 0.500
untrained 0.6025 0.4 0.5
SFT-only 0.7 0.575 0.4
GRPO 0.700 0.575 0.475
PPO 0.6625 0.700 0.55
CISPO 0.6875 0.4875 0.4
VSPO 0.7125 0.725 0.55
Table 1: The performance results of VSPO with baselines in long-form QA.
Short-Form QA
Results are shown in Table2. We observe that: (1) Across all seven datasets, VSPO+PRS achieves the highest average score (0.419), outperforming all other RL baselines, including PPO+PRS and GRPO+PRS. This demonstrates the effectiveness of combining VSPO with the proposed progressive reward shaping in improving short-form reasoning tasks for LLM agents. (2) Comparing reward designs, PRS consistently improves performance over the rule-based exact match reward (EM) across PPO, GRPO, and VSPO, confirming that reward shaping can better guide long-horizon reasoning even in short-form question answering. (3) Overall, VSPO yields stable gains in most datasets without sacrificing performance on any, indicating strong robustness across both in-domain (NQ, HotpotQA) and out-of-domain (TriviaQA, PopQA, 2wiki, Musique and Bamboogle) benchmarks.
NQ TriviaQA PopQA HotpotQA 2wiki Musique Bamboogle Avg
Direct Inference 0.134 0.408 0.14 0.183 0.25 0.031 0.12 0.181
SFT 0.318 0.354 0.121 0.217 0.259 0.066 0.112 0.207
PPO+EM(SEARCH-R1) 0.393 0.610 0.397 0.370 0.414 0.146 0.368 0.385
GRPO+EM(SEARCH-R1) 0.429 0.623 0.427 0.386 0.346 0.162 0.400 0.396
VSPO+EM 0.433 0.623 0.425 0.396 0.350 0.162 0.390 0.397
PPO+PRS 0.410 0.610 0.400 0.386 0.410 0.157 0.400 0.396
GRPO+PRS 0.440 0.639 0.420 0.400 0.39 0.171 0.413 0.410
VSPO+PRS 0.440 0.645 0.416 0.408 0.410 0.171 0.435 0.419
Table 2: The performance results of VSPO and PRS with baselines in shor-form QA.
@@
We analyze some issues based on long-form QA, and the same applies to the case of short-form QA.
@@
The training reward and eval reward of both RL algorithm are show in Firure 2.
@@
Figure 2: Reward curves on the training and validation sets for RL algorithm in long-form QA.
@@
VSPO achieved the best performance based on rewards on the training and validation sets. As the number of training steps increases, PPO gradually converges to the same level as VSPO and GRPO on training set.
@@
Among all RL algorithms, VSPO was the first to receive the best validation reward, this indicates that filling in samples that cannot be helpful for learning through value-based sampling makes the model more focused on high-value samples, which helps to increase convergence speed and achieve better performance.
The Effective of PRS
To demonstrate the effectiveness of PRS, we use rule-baesd 0-1 reward as the baseline, using VSPO as the RL algorithm. This 0-1 reward use Qwen2.5-72B-Instruct as the verifier, determine whether the model output passes based on the annotated answers, if passed, the reward is 1, otherwise it is 0.
The training reward and eval reward of both reward design are show in Figure 3. Note that the scales of the two rewards are inconsistent, so we normalize PRS using the maximum and minimum values within the range of PRS. From curves on training set, we can see that, compared to the 0-1 reward, PRS achieved a relatively good convergence level earlier. And the reward of PRS on the validation set is also higher than 0-1 reward.
Figure 3: Reward curves on the training and validation sets for two reward designs.
The result on the test set is shown in Table 3. From this table we can see that PRS outperforms 0-1 reward on all query types, which proves the effectiveness of PRS.
$Q_{simple}$ $Q_{multiq}$ $Q_{multim}$
0-1 reward 0.6905 0.600 0.500
PRS 0.7125 0.725 0.550
Table 3: Results on the test set of 0-1 reward and PRS.
@@
Figure 4 shows the entropy loss of VSPO when using 0-1 reward and PRS as training reward. Entropy loss reflects the uncertainty of policy model in selecting the current action. We observe that PRS reduces uncertainty to a lower level earlier than the binary reward and achieves better test set performance. This is not entropy collapse (which would manifest as degraded test metrics), but rather efficient convergence. Therefore, we infer that PRS, as a curriculum-learning-inspired reward design, effectively guides the policy’s exploration and enables faster convergence to high-quality strategies. In contrast, the binary reward leads to inefficient exploration under sustained high uncertainty.
@@
Ablation Analysis of VSPO
@@
To evaluate the effectiveness of value-based sample and value smoothing clipping of VSPO, we conduct experiments using no sample without clip (equivalent to GRPO), random sample without clip, value-based sample without clip, random sample with clip and value-based sample with clip (equivalent to VSPO), respectively. The result on the test set is shown in Table 4. The table demonstrates that removing either value-based sampling or value smoothing clipping significantly degrades performance, confirming the necessity of both components. Notably, any sampling strategy without clipping leads to training instability and poor performance, while value-based sampling with clipping achieves the best results across all query types.
$Q_{simple}$ $Q_{multiq}$ $Q_{multim}$
w/o sample, w/o clip
0.700 0.575 0.475
random sample, w/o clip
0.460 0.160 0.100
value-based sample, w/o clip
0.450 0.175 0.300
random sample, w/ clip
0.500 0.300 0.025
value-based sample, w/ clip
0.7125 0.725 0.550
@@
Table 4: Results on the test set of different ablation settings.
@@
We further analyze the impact of value-based sampling and value smoothing clipping through KL divergence, training rewards, and validation rewards. Figure 5 presents these metrics across different configurations. Two key observations emerge:
(1) Value smoothing clipping stabilizes training. Without clipping, both random sampling and value-based sampling exhibit abnormally high KL divergence, leading to extremely unstable training. In contrast, configurations with clipping maintain stable KL divergence throughout training, demonstrating that value smoothing clipping is essential for training stability.
(2) Value-based sampling is necessary for performance gains. From the training and validation reward curves, we observe that random sampling fails to improve performance even when combined with value smoothing clipping. However, value-based sampling with clipping achieves substantial improvements on both training and validation sets, confirming the necessity of intelligent sample selection based on learning value.
@@
(b) Reward on eval set
(c) Reward on train set
Figure 5: The KL Loss and Reward Curve.

### v2

@@ A Dynamic Advantage Weight Perspective of VSPO
VSPO can be interpreted as task-value-based reweighting rather than random sampling. Consider the general policy gradient (Sutton et al. 1999) with advantage function $A_{t}$:
@@
To evaluate the effectiveness of PRS and VSPO in training LLM-based TIR agent, we conduct experiments on the following two types of long-horizon reasoning tasks: short-form QA and long-form QA. In these tasks, the agent has access to at least one information retrieval tool to answer the question.
In short-form QA, we use Qwen2.5-3B-Instruct (Yang et al. 2024) and in long-form QA, we use the Qwen3-14B (Team 2025) as the training model. All of them uses full parameter training.
We adopt Trinity-RFT (Pan et al. 2025) as our training framework.
For VSPO, we set $T=0.1$ and $\alpha=2$. The rollout num of VSPO, CISPO and GRPO is 5. The learning rate of all algorithms is 1e-7. Training is performed on a single node with 8 H20 GPUs (141GB). To optimize GPU memory usage, we enable gradient checkpointing and use Fully Sharded Data Parallel (FSDP) with CPU offloading.
@@
The number of training and testing sets is 1678 and 334, respectively. We manually annotate reference answers based on our internal knowledge base.
For expert data used in SFT, we use knowledge distillation from the Qwen3-235B-A22B (Team 2025) to construct high-quality trajectory data. Specifically, we collect online service logs from Qwen3-235B-A22B and apply rule-based filtering and LLM-as-a-Judge cleaning to produce expert demonstrations.
@@
Following (Jin et al. 2025), we use seven benchmark datasets: (1) General Question Answering: NQ (Kwiatkowski et al. 2019), TriviaQA (Joshi et al. 2017), and PopQA (Mallen et al. 2023). (2) Multi-Hop Question Answering: HotpotQA (Yang et al. 2018), 2WikiMultiHopQA (Ho et al. 2020), Musique (Trivedi et al. 2022), and Bamboogle (Press et al. 2023) as Short-Form QA train and test data.
@@
We use GPT-4.1 as an LLM judge to determine whether the responses contains hallucinations and whether it matches the annotated reference answers. The evaluation prompts are provided in Appendix B.
@@
We use Exact Match (EM) as the evaluation metric, computed against the golden answers provided by each benchmark dataset.
@@
To verify the effectiveness of PRS, we compare it with rule-based 0-1 reward to demonstrate the effectiveness of PRS. All reward functions use GRPO as the RL algorithm. For long-form QA, the 0-1 reward use Qwen3-235B-A22B-Instruct-2507 (Team 2025) as the verifier, determine whether the model output passes based on the annotated reference answers, if passed, the reward is 1, otherwise it is 0. For short-form QA, using exact match as rule-based 0-1 reward.
To verify the effectiveness of VSPO, we compare it with GRPO, PPO (Schulman et al. 2017a) and SFT-only. All RL algorithms use the same PRS reward.
@@
The main results of PRS
The main results comparing PRS with 0-1 reward are presented in Table 1 for long-form QA and Table 2 for short-form QA. Under the same RL algorithm, we can found that compared to 0-1 reward, PRS achieved better results. In long-form OA, $PRS_{long}$ has increased reference answer pass rate by $0.62\%$ and hallucination Pass Rate by $1.7\%$ compared to 0-1 reward. In short-form OA, $PRS_{short}$ has increased exact match score by $2.74\%$ compared to 0-1 reward.
Qwen3-14B
Methods $Q_{simple}$ $Q_{multiq}$ $Q_{multim}$ Avg.
Reference Answer Pass Rate
Untrained 0.5513 0.6999 0.5000 0.5897
GRPO + 0-1 reward 0.6955 0.8017 0.6000 0.7255
GRPO + $PRS_{long}$ 0.7144 0.8183 0.7150 0.7300
Hallucination Pass Rate
Untrained 0.7821 0.8167 0.7143 0.7893
GRPO + 0-1 reward 0.8782 0.9333 0.8600 0.8925
GRPO + $PRS_{long}$ 0.8982 0.9333 0.9000 0.9077
Table 1: Results on the test set of 0-1 reward and PRS for long-form QA.
Qwen2.5-3B-Instruct
Methods NQ TriviaQA PopQA HotpotQA 2wiki Musique Bamboogle Avg.
Exact Match (EM)
Untrained 0.0415 0.1751 0.0849 0.0622 0.0891 0.0082 0.1667 0.0960
GRPO + 0-1 reward 0.3213 0.5022 0.3457 0.3000 0.2586 0.1120 0.0833 0.3390
GRPO + $PRS_{short}$ 0.3518 0.4987 0.3562 0.3100 0.2600 0.1245 0.3333 0.3483
Table 2: Results on the test set of 0-1 reward and PRS for short-form QA.
The main results of VSPO
The main results comparing VSPO with baseline methods across are presented in Table 3 for long-form QA and Table 4 for short-form QA. Under the same reward function, we can found that compared to other RL algorithm, VSPO achieved better results. In long-form QA, VSPO has increased reference answer pass rate by $5.34\%$ compared to GRPO. In short-form OA, $PRS_{short}$ has increased exact match score by $3.16\%$ compared to GRPO.
Qwen3-14B
Methods $Q_{simple}$ $Q_{multiq}$ $Q_{multim}$ Avg.
Reference Answer Pass Rate
Untrained 0.5513 0.6999 0.5000 0.5897
SFT 0.6405 0.7500 0.6000 0.6687
GRPO + $PRS_{long}$ 0.7144 0.8183 0.7150 0.7300
PPO + $PRS_{long}$ 0.700 0.8213 0.7020 0.7327
VSPO + $PRS_{long}$ 0.7147 0.9000 0.8557 0.7690
Hallucination Pass Rate
Untrained 0.7821 0.8167 0.7143 0.7893
SFT 0.8079 0.8167 0.7556 0.8086
GRPO + $PRS_{long}$ 0.8982 0.9333 0.9000 0.9077
PPO + $PRS_{long}$ 0.8282 0.8633 0.9000 0.8399
VSPO + $PRS_{long}$ 0.8662 0.9000 0.8571 0.8750
Table 3: Results on the test set of RL algorithm for long-form QA.
Qwen2.5-3B-Instruct
Methods NQ TriviaQA PopQA HotpotQA 2wiki Musique Bamboogle Avg.
Exact Match (EM)
Untrained 0.0415 0.1751 0.0849 0.0622 0.0891 0.0082 0.1667 0.0960
SFT 0.2490 0.2920 0.1040 0.1860 0.2480 0.0440 0.1120 0.1760
GRPO + $PRS_{short}$ 0.3518 0.4987 0.3562 0.3100 0.2600 0.1245 0.3333 0.3483
PPO + $PRS_{short}$ 0.3210 0.4887 0.3562 0.3200 0.2650 0.1200 0.3325 0.3333
VSPO + $PRS_{short}$ 0.3600 0.5000 0.3634 0.3100 0.2650 0.1367 0.3489 0.3593
Table 4: Results on the test set of RL algorithm for short-form QA.
@@
We analyze some issues based on long-form QA and short-form QA.
@@
The training reward and test set metric of both RL algorithm are show in Firure 2.
@@
Figure 2: Reward curves on the training and test set metric(Reference Answer Pass Rate) for RL algorithm in long-form QA.
@@
VSPO achieved the best performance based on rewards on the train set and reference answer pass rate on the test sets. As the number of training steps increases, PPO gradually converges to the same level as VSPO and GRPO on training set.
@@
Among all RL algorithms, VSPO was the first to receive the best test metric performance, this indicates that filling in tasks that cannot be helpful for learning through value-based sampling makes the model more focused on high-value tasks, which helps to increase convergence speed and achieve better performance.
The Effect of $T$ and $\alpha$ of VSPO
We analyze the effect of $T$ and $\alpha$ of VSPO. Based on equation 19, We can analyze that $T$ controlling the sharpness of value distribution. If $T$ tends towards 0, the value-based sampling will be concentrated in high learning value tasks. If $T$ tends towards positive infinity, the value-based sampling will become flat and uniform and weaken the gap in learning value of tasks.
$\alpha$ controlls the maximum amplitude of sample update. The larger the value of $\alpha$, the more radical updates are allowed, while the smaller the value, the more conservative updates are allowed.
The results of different $\alpha$ and $T$ are shown in Table 5 and the reference answer pass rate curve of test set is shown in Figure 3.
Qwen3-14B
Methods $Q_{simple}$ $Q_{multiq}$ $Q_{multim}$ Avg.
Reference Answer Pass Rate
Untrained 0.5513 0.6999 0.5000 0.5897
$\alpha=2$, $T=1$ 0.7019 0.8833 0.8571 0.7556
$\alpha=2$, $T=0.5$ 0.7276 0.8250 0.8571 0.7579
$\alpha=2$, $T=0.1$ 0.7147 0.9000 0.8557 0.7690
$\alpha=4$, $T=0.5$ 0.6923 0.875 0.8571 0.7134
$\alpha=6$, $T=0.5$ 0.7051 0.8750 0.8571 0.7218
no clip, $T=0.5$ 0.6955 0.8917 0.996 0.7197
Hallucination Pass Rate
Untrained 0.7821 0.8167 0.7143 0.7893
$\alpha=2$, $T=1$ 0.8782 0.8667 0.8658 0.8747
$\alpha=2$, $T=0.5$ 0.8846 0.8167 0.8571 0.8655
$\alpha=2$, $T=0.1$ 0.8662 0.9000 0.8571 0.8750
$\alpha=4$, $T=0.5$ 0.8782 0.9167 0.7143 0.8494
$\alpha=6$, $T=0.5$ 0.8526 0.8667 0.4286 0.8177
no clip, $T=0.5$ 0.8526 0.8333 0.8714 0.8075
Table 5: Results on the test set of different $\alpha$ and $T$ of VSPO for long-form QA.
Figure 3: Reference Answer Pass Rate of test set of different $\alpha$ and $T$.
From Table 5 and Figure 3, We can analyze that, in VSPO:
  • When $\alpha$ is fixed, a lower $T$ will bring better results and as $T$ decreases, the best result will be achieved earlier. This is because a lower $T$ allows value-based sampling to be repeated more times on more valuable tasks, which allows more valuable tasks to have a relatively larger gradient modulus to update themselves.
  • When $T$ is fixed, raising $\alpha$ did not bring any improvement in effectiveness and $\alpha=+\infty$ (means no value smoothing clipping, corresponding to no clip) lead to a decrease in the effect. This is because the main function of $\alpha$ is to limit the changes in advantages in a smooth manner. If the magnitude of $\alpha$’s changes is too large, it may lead to unstable updates and reduce the effectiveness of training.
The Effect of PRS
Not limited to the metric results shown in Table 1 and Table 2 on the test set, we will further analyze the effect of PRS.
PRS can reduce the proportion of zero advantage tasks in batch.
Figure 4 shows the proportion of zero advantage tasks at each training step when using 0-1 reward and $PRS_{long}$ (The RL algorithm is GRPO).
We can see that, when using 0-1 reward, about 60% of tasks for batch in each step do not bring any learning signals. For PRS, based on the summation of various components, it can generate rich partial order, so the proportion of zero advantage tasks is greatly reduced.
Figure 4: Proportion of zero advantage tasks when training when using 0-1 reward and PRS.
PRS can help policy explore strategies more efficiently.
@@
Figure 5 shows the entropy loss of VSPO when using 0-1 reward and PRS as training reward. Entropy loss reflects the uncertainty of policy model in selecting the current action. We observe that PRS reduces uncertainty to a lower level earlier than the 0-1 reward and achieves better test set performance. This is not entropy collapse (which would manifest as degraded test metrics), but rather efficient convergence. Therefore, we infer that PRS, as a curriculum-learning-inspired reward design, effectively guides the policy’s exploration and enables faster convergence to high-quality strategies. In contrast, the 0-1 reward leads to inefficient exploration under sustained high uncertainty.
@@
PRS has better improvement on weaker models
Adding Qwen3-8B to long-form QA and Qwen2.5-7B-Instruct to short-form QA, the results comparing PRS with 0-1 reward are presented in Table 6 for long-form QA and Table 7 for short-form QA.
In long-form OA, $PRS_{long}$ has increased reference answer pass rate by $1.27\%$ ($0.62\%$ for Qwen3-14b) and hallucination Pass Rate by $22.01\%$ ($1.7\%$ for Qwen3-14b) compared to 0-1 reward. In short-form QA, $PRS_{short}$ has increased exact match score by $2.74\%$ compared to 0-1 reward ($1.08\%$ for Qwen2.5-7b-Instruct). It can be seen that PRS brings more improvements to weaker models. We believe that the reason is that weak models have weaker exploration ability in the process of RL compared to strong models, so the lack of a guiding reward function (such as 0-1 reward) is a difficult learning signal for weak models. On the contrary, PRS based on curriculum learning gradually decomposes complex tasks through reward settings, which helps weak models gradually explore and improve strategy performance.
Qwen3-8B
Methods $Q_{simple}$ $Q_{multiq}$ $Q_{multim}$ Avg.
Reference Answer Pass Rate
Untrained 0.5353 0.1083 0.3571 0.4148
GRPO + 0-1 reward 0.6731 0.1000 0.6429 0.5180
GRPO + $PRS_{long}$ 0.6506 0.1917 0.5714 0.5246
Hallucination Pass Rate
Untrained 0.7372 0.1000 0.7143 0.5650
GRPO + 0-1 reward 0.8974 0.1167 0.8971 0.6873
GRPO + PRS 0.8654 0.7667 0.8571 0.8386
Table 6: Results on the test set of 0-1 reward and $PRS_{long}$ for long-form QA.
Qwen2.5-7B-Instruct
Methods NQ TriviaQA PopQA HotpotQA 2wiki Musique Bamboogle Avg.
Exact Match (EM)
Untrained 0.2770 0.4872 0.3289 0.2797 0.2411 0.1037 0.2320 0.3210
GRPO + 0-1 reward 0.3500 0.5438 0.3754 0.3600 0.3095 0.1286 0.2996 0.3855
GRPO + $PRS_{short}$ 0.3573 0.5623 0.3815 0.3757 0.3174 0.0996 0.3333 0.3897
Table 7: Results on the test set of 0-1 reward and $PRS_{short}$ for short-form QA.
Ablation Analysis of Value-based Sampling and Value Smoothing Clipping
@@
To evaluate the effectiveness of value-based sampling and value smoothing clipping of VSPO, we conduct experiments using no sample without clip (equivalent to GRPO), random sample without clip, value-based sample without clip, random sample with clip and value-based sample with clip (equivalent to VSPO), respectively. The result on the test set is shown in Table 8. Table 8 demonstrates that removing either value-based sampling or value smoothing clipping significantly degrades performance, confirming the necessity of both components. Notably, any sampling strategy without clipping leads to a decrease in performance, while value-based sampling with clipping achieves the best performance.
Qwen3-14B
Methods $Q_{simple}$ $Q_{multiq}$ $Q_{multim}$ Avg.
Reference Answer Pass Rate
w/o sample w/o clip (GRPO) 0.7144 0.8183 0.7150 0.7300
random sample w/o clip 0.4120 0.1600 0.5400 0.3100
value-based sample w/o clip 0.6955 0.8917 0.9960 0.7197
random sample w clip 0.6322 0.5790 0.7150 0.6335
value-based sample w clip (VSPO) 0.7147 0.9000 0.8557 0.7690
Hallucination Pass Rate
w/o sample w/o clip (GRPO) 0.8982 0.9333 0.9000 0.9077
random sample w/o clip 0.5368 0.2300 0.5734 0.4149
value-based sample w/o clip 0.8526 0.8333 0.8714 0.8075
random sample w clip 0.7950 0.6333 0.8876 0.8256
value-based sample w clip (VSPO) 0.8662 0.9000 0.8571 0.8750
@@
Table 8: Results on the test set of different ablation settings for long-form QA.
@@
We further analyze the impact of value-based sampling and value smoothing clipping through KL divergence on train set and reference answer pass rate on test set. Figure 6 presents these metrics across different configurations. Two key observations emerge:
(1) Value smoothing clipping stabilizes training. When using random sampling, without clipping causes random sampling abnormally high KL divergence and low metric on test set, which leading to extremely unstable training. With clip can change the low performance of random sampling to some extent. When using value-based sampling, without clipping will have a larger KL loss and lower reference answer pass rate on test set than with clip. The above information can verify the effectiveness of value smoothing clipping. By limiting the changes in advantages in a smooth manner, value smoothing clipping stabilizes training while improving sampling performance.
(2) Value-based sampling is necessary for performance gains. When using value smoothing clipping, value-based sampling have a higher reference answer pass rate on test set on random sampling. When not using value smoothing clipping, random sampling have abnormally high KL divergence and lower metric on test set than value-based sampling. This confirming the necessity of intelligent sample selection based on learning value.
@@
(b) Reference Answer Pass Rate on test set
Figure 6: The KL Loss and Reference Answer Pass Rate on test set.